OpenStudy (anonymous):

Last year the Mavericks ratio of wins to losses was 4:3, of the 84 games they played, how many did they win?

OpenStudy (anonymous):

Well, 4+3 = 7: Then 84/7 = 12 Then you do 4*12 for wins and 3*12 = losses.